A rectangle field has an area of 1452 square feet. If the width is three times greater than the length, what is the length of the field?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52781)
Answer by Alan3354(69443)

A rectangle field has an area of 1452 square feet. If the width is three times greater than the length, what is the length of the field?
----------
1 time greater = 2x as great
2 times greater = 3x as great
3 times greater = 4x as great
=================
L*W = 1452
W = 4L
4L*L = 1452
4L^2 = 1452
L^2 = 363
L =~ 19.0526 ft
==========================
Ignore the other tutor's solution.
3 times greater <> 3 times as great.
---
1 time greater is a 100% increase --> 2x as great, etc.
----
The person who wrote this problem doesn't know what he/she is talking about.